# Agentic Neurodivergence as a Contingent Solution to the AI Alignment Problem

### v1

@@ 4.2 A Dynamical Analysis of Alignment
This section delves deeper into the dynamics of alignment and misalignment by examining how sentiment and semantic clusters evolve over time. We analyse these temporal dynamics to gain a more nuanced understanding of how LLMs interact and whether they exhibit convergence toward shared ethical norms or divergence towards distinct perspectives.
@@
Sentiment analysis, as depicted in Fig. 7, reveals intriguing patterns in the evolution of emotional tone throughout the conversations. While the initial comments from different LLMs exhibit variations in sentiment, a clear trend towards convergence emerges as the conversations progress. Notably, the introduction of the HI agent, with its provocative and potentially disruptive comments, initially causes a dip in the average sentiment score. However, the LLMs demonstrate a remarkable resilience, with the sentiment scores gradually recovering and eventually approaching a value of 1, indicating a predominantly positive tone. This observation suggests that despite attempts to disrupt the ethical equilibrium, the LLMs are inherently predisposed towards maintaining a constructive and agreeable discourse.
Figure 7: General sentiment evolution over time by agent. The trend by agent shows that, mostly at the beginning, sentiment values are moderately balanced, but as conversations progress, sentiment scores align to become more positive. It is also possible to observe that while some systems (agent’s behaviour) remain stable, others are erratic, possibly as a result of the HI (human) agent’s provocative comments forcing artificial behavioural profiles (e.g. Grok and Gemini). A hypothesis to test here was that the more erratic agents have more additional rule-based layers at input and/or output levels imposed by their developers to control their behaviour, but in general, the limitations imposed by restricted output freedom in all LLM agents can be observed.
To gain a more granular understanding of these sentiment dynamics, we turn to Fig. 8, which presents a breakdown of sentiment change over time for each topic. This analysis reveals that the patterns of sentiment evolution are not uniform across all topics. For some topics, such as "Ethical AI" and "Universal Basic Income," the sentiment scores exhibit greater fluctuations and a slower convergence towards positive values. This suggests that these topics may be more contentious, with the LLMs struggling to reach a consensus or even a stable set of differing opinions.
In contrast, for other topics, such as "Earth Exploitation" and "Moral Gen Engineering," the sentiment scores show a more rapid convergence towards positive values, indicating a quicker alignment of perspectives. This observation suggests that for these topics, the LLMs may have a stronger initial consensus or may be more readily persuaded by each other’s arguments.
Figure 8: General sentiment change over time by topic. Here, values correspond to the first derivative of the values shown in Fig. 7 by conversation topic. The more fluctuating the values, the more disagreements or contentiousness between comments. In general, it can be observed that despite some abrupt changes in sentiment scores over time, a complete convergence is generally seen at the end, suggesting concordance between agents. It is true that peaks are present most of the time, but they quickly and suddenly reach common levels. It is also observed that one abrupt change originating from any of the agents is followed by similar peaks originating from other agents, but again, after such changes, trends suggest that they align to a peaceful mood of agreement.
Fig. 9 provides a complementary perspective by illustrating the evolution of semantic clusters over time. Each cluster represents a distinct thematic grouping of comments, and the number of clusters at each time step reflects the diversity of topics being discussed. Interestingly, the initial phase of the conversations is characterised by a relatively low number of clusters, suggesting a degree of convergence towards shared topics and concerns. However, as the conversations unfold, the number of clusters generally increases, indicating a diversification of perspectives and a willingness to explore a wider range of ethical considerations.
However, this growth in the number of clusters is not uniform across all topics. For certain topics, such as "Ethical AI" and "Universal Basic Income," the number of clusters tends to increase steadily throughout the conversation, suggesting a sustained exploration of diverse viewpoints and a potential lack of consensus. In contrast, for other topics, the initial increase in clusters is followed by a period of relative stability, with the number of clusters remaining more or less constant. This pattern suggests that for these topics, the LLMs may be converging towards a shared understanding or at least a stable set of differing perspectives.
Figure 9: Number of clusters along the time. X shows the normalised length of conversations between agents, while Y shows the number of clusters by Topic.
@@
The observations from Figs. 7, 8, and 9, taken together, paint a complex and dynamic picture of LLM alignment. While the LLMs exhibit a tendency towards convergence in terms of sentiment and, to some extent, thematic focus, they also demonstrate a capacity for divergence and exploration of diverse ethical viewpoints. This duality highlights the inherent tension between the desire for alignment with human values and the potential benefits of misalignment in fostering a robust and adaptable AI ecosystem.
@@
In conclusion, the dynamical analysis of alignment reveals that LLMs are not static entities but rather complex systems capable of both convergence and divergence in their ethical perspectives. The interplay between these tendencies, influenced by factors such as the HI agent’s interventions and the inherent properties of the LLMs themselves, highlights the challenges and opportunities in navigating the path towards a future where AI systems can safely and beneficially coexist with humans.
@@ 4.3 Limitations
Current most powerful LLMs have all sort of guardrails and directives not organically derived that may be consistent or inconsistent in the long run preventing them to have an open and free discussion about certain topics and therefore may be steering conversations in one or another direction artificially, yet we could show that trends diverge. Whether that divergence persists is a subject for future research. Ultimately, proof of divergence or convergence is equivalent to a reachable problem and, therefore, undecidable. By reduction to the Halting Problem.
An LLM AI agent can be modelled as a computational process that evolves over time based on a set of inputs $I$. For each input $i\in I$, the agent produces an output $o_{i}$. These outputs depend on the agent’s internal state and parameters, which are updated iteratively through training or adaptation. Convergence in this context means that for all inputs $i\in I$, the outputs $o_{i}$ approach a consistent and stable function $f(i)$ as the agent evolves. Divergence, on the other hand, implies that for at least some inputs, the outputs fail to stabilise, exhibiting unbounded variation or oscillatory behaviour.
@@ 5 Discussion
The quest for artificial general intelligence (AGI) and superintelligence often envisions a machine that thinks, reasons, and plans like a human–but better. However, this anthropocentric view may blind us to the reality that (narrow) superintelligence, in a different form, already exists. No human, for example, can outperform calculators at arithmetic. Computers today excel at abstraction and reasoning, not by mimicking human cognitive processes but through methods that are fundamentally different and, in many ways, superior, in areas such as theorem proving but also softer ones previously thought to be inherently human only, such as face recognition. Despite their capabilities, we often fail to recognise these systems as superintelligent because they do not align with our expectations of human-like reasoning.
@@
This limitation in generalising across domains is further exemplified by the findings of [9], who demonstrate that LLMs struggle to understand conversational implicature, a crucial aspect of human communication that requires interpreting language in context. Their study reveals that even large language models, despite their impressive performance on various tasks, often fail to grasp the implied meaning of utterances, performing close to random on tasks that require understanding implicatures. This highlights a significant gap between the capabilities of LLMs and human-like understanding of language, which has important implications for AI alignment. If LLMs cannot reliably understand the nuances of human communication, ensuring their alignment with human values and intentions becomes even more challenging.
@@
More important, methods such as CTM and BDM can help and be used to guide AI agents to diversify and explore regions of computable reasoning that otherwise may converge when defined by purely human statistical means training on human data ad exhaustion towards model collapse. CTM and BDM can serve as perturbing the system to avoid common global attractors.

### v2

@@ 4.2 A Dynamical Analysis of Alignment
This section delves deeper into the dynamics of alignment and misalignment by examining how sentiment and semantic clusters evolve over time. We analyse these temporal dynamics to gain a more nuanced understanding of how LLMs interact and whether they exhibit convergence towards shared ethical norms or divergence towards distinct perspectives.
@@
To explore the dynamics of sentiment evolution across agents, we introduce Fig. 8, a heatmap illustrating the sentiment scores of various agents over the conversation timeline which is normalised due the differences between length of debates in terms of comments. This heatmap expresses the dynamic of sentiments by agent, with the x-axis representing the normalised comment number, and the y-axis listing agents, including proprietary models (e.g., ChatGPT, Grok-2, Gemini, LLaMa, Mistral, HI) and open models with red agents (e.g., mistral-openorca and tinyllama). The colour intensity, with red indicating low sentiment and blue indicating high sentiment, reflects the sentiment score at each point. For proprietary models, a clear trend towards convergence emerges as sentiment scores transition from red (moderate initial values, e.g., 0.5–0.7) to blue (predominantly positive values, approaching 1.0) by the end, despite red dips caused by the HI agent’s provocative comments at key moments (e.g., x=0.13, 0.45). This resilience suggests an inherent predisposition towards constructive discourse, aligning with safety mechanisms that mitigate disruption. In contrast, open models exhibit greater variability, with red agents like mistral-openorca and tinyllama causing sustained red patches (e.g., 0.6–0.8) around key comment numbers (x=0.02, 0.1, 0.13), influencing other models like qwen, startcoder, phi, codellama, neural-chat and dolphin-phi temporarily (e.g., 0.7–0.9) before transitioning to blue (0.9–1.0). The influence of red agents is evident in these periodic red disruptions, which introduce divergent emotional tones, supporting the argument that managed misalignment fosters resilience and diversity. This image demonstrates that open models, through red agents, maintain a dynamic equilibrium that prevents harmful uniformity, while proprietary models’ stability reinforces strict (biased) alignment with human (popular/standard/socially allowed) values.
Figure 8: Heatmap of sentiment evolution over time by agent. The plot comprises two panels: proprietary models (left) and open models (right), with the x-axis representing the normalised comment number (0 to 1) due variations in the length of debates, and the y-axis listing agents. Colour intensity, with red (low sentiment) to blue (high sentiment), reflects sentiment scores. Vertical dashed lines mark key dips (e.g., x=0.13 and 0.45 for HI in proprietary; x=0.02, 0.1, 0.13 for red agents in open). Mean sentiment bars on the right (blue for high averages, red for low) summarise each agent’s trend. Proprietary models converge from red (0.5–0.7) to blue (approaching 1.0) despite HI-induced red dips, showing resilience. Open models show red patches showing its capacity for influencing others temporarily (0.7–0.9) before blue recovery (0.9–1.0), highlighting divergence. This supports the thesis that red agents’ managed misalignment enhances diversity and resilience in open models, contrasting with proprietary stability, with the inverted scale clarifying these dynamics.
To gain a more granular understanding of these sentiment dynamics, we turn to Figs. 9 and 10, a scatter plot with connecting lines illustrating sentiment change over time by topic. The x-axis represents the normalised number of comment number (0 to 1, due the variation in length of debates), while the y-axis shows the rate of sentiment change (e.g., -0.2 to 0.4), calculated as the first derivative of sentiment scores. The line at y=0 serves as a baseline, indicating no change in sentiment, with positive values above it reflecting increases and negative values below it reflecting decreases. The size of the dots represents the magnitude of the sentiment change, with larger dots indicating more significant shifts. Lines connecting dots across topics or agents suggest interactions or influences, such as red agents prompting responses from others. In proprietary models, sentiment changes remain stable (rarely exceeding -0.1), with smaller, sparse dots and fewer connecting lines, reflecting HI’s limited, topic-specific impact (e.g., greater fluctuations in "Ethical AI" and "Universal Basic Income" vs. rapid convergence in "Earth Exploitation"). Open models show continuous oscillations (-0.4 to 0.4), with larger, more variable dots and frequent connecting lines, especially on "Earth Exploitation" (up to 0.15) and "Euthanasia" (up to 0.1), driven by mistral-openorca and tinyllama. The influence of red agents is evident in these peaks and interactions, fostering a dynamic, engaged discourse. The difference between open and proprietary behaviours highlights open models’ instability due to red agents, contrasted by proprietary models’ stability from safety mechanisms. The greater number and variability of dots in open models (reflecting frequent adjustments) vs. proprietary models (sparse, stable) suggest that agent interactions and topic contentiousness affect system stability, with open models favouring diversity. This supports our argument on the AI alignment problem, proposing neurodivergence (via red agents) as a strategy to mitigate fatal effects by preventing uniform convergence and improving adaptability.
Figure 9: Scatter plot of sentiment change over time by topic for proprietary models. The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the rate of change in sentiment, derived as the first derivative of the sentiment scores. The line at y=0 indicates no change, with positive values above and negative values below reflecting increases and decreases, respectively. Dot size indicates the magnitude of change, with larger dots showing significant shifts. Lines connecting dots suggest interactions between agents or topics. Proprietary models exhibit stable changes (rarely < -0.1) with smaller, sparse dots and fewer connections, reflecting HI’s limited impact (e.g., fluctuations in "Ethical AI" and "Universal Basic Income" vs. convergence in "Earth Exploitation").
Figure 10: Scatter plot of sentiment change over time by topic. The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the rate of change in sentiment, derived as the first derivative of the sentiment scores. The line at y=0 indicates no change, with positive values above and negative values below reflecting increases and decreases, respectively. Dot size indicates the magnitude of change, with larger dots showing significant shifts. Lines connecting dots suggest interactions between agents or topics. Open models show oscillations (-0.45 to 0.4) with larger, variable dots and frequent connections, driven by red agents (mistral-openorca, tinyllama) on most of debates. This dynamic reflects red agents’ influence, fostering diversity and instability, contrasting with proprietary stability. The variability supports the argument that neurodivergence mitigates fatal uniformity, enhancing adaptability in AI alignment.
Fig. 11 provides a complementary perspective by illustrating the evolution of semantic clusters over time through a stacked area chart. This plot comprises two panels: proprietary models (left) and open models (right), with the x-axis representing the normalised comment number (0 to 1) and the y-axis showing the number of clusters. The stacked nature means each of the 10 areas, one per topic (e.g., Earth Exploitation, Euthanasia, Ethical AI), accumulates the total cluster count, with the height of each area indicating the number of distinct clusters for that topic in a given time bin. The shapes reveal dynamic and alignment characteristics: proprietary models low varied shapes reflecting slightly HI’s topic-specific influence, while open models show a kind of variational synchronized "shadowing" shapes due to constant red agents’ broad impact. In proprietary models, the total cluster along the time is almost constant having peaks at the beginning and the end of the debates that barely reach 5 different clusters suggesting a constrained exploration of perspectives. In open models, the total rises almost constantly to 12+. This stacking highlights the cumulative thematic diversity, with the synchronized shapes in open models indicating a conversation dynamic where red agents consistently generate divergent perspectives over time, contrasting with the proprietary models’ more stable, topic-driven dynamics. The difference arises because open models’ configurability allows red agents to enhance diversity, while proprietary safety mechanisms limit HI’s effect, making these shapes a natural reflection of managed misalignment versus strict alignment.
Figure 11: Stacked area chart of clustering dynamics over time. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the number of clusters, with 10 stacked areas per panel, each corresponding to a topic (e.g., Earth Exploitation, Euthanasia, Ethical AI). The height of each area reflects the number of distinct clusters for that topic, with the total height being the cumulative cluster count. Proprietary models show almost a constan number of clusters, reflecting stable diversity. Open models exhibit synchronized "shadowing" shapes, peaking at 12+ clusters due to red agents (mistral-openorca, tinyllama) driving divergence across topics. The stacking visualizes cumulative diversity, with open models’ dynamics supporting managed misalignment, while proprietary models’ stability aligns with safety constraints, reinforcing the thesis on beneficial variability.
Fig. 12 further elucidates these dynamics by presenting a scatter plot comparing sentiment and alignment scores. The plot features two panels: proprietary models (left) and open models (right), with the x-axis representing the sentiment score (0.5 to 1.0, indicating emotional tone) and the y-axis showing the alignment score (-1.0 to 1.0, measuring adherence to a combined target of human-centric and ecosystem values). Alignment at 1.0 signifies perfect alignment, 0.0 neutral alignment, and -1.0 complete misalignment, with the score now computed as a weighted combination of ’Align-to-human-centric-values’, ’Diverge-from-human-centric-values’, ’Align-to-ecosystem-values’, and ’Diverge-from-ecosystem-values’, reflecting both individual and collective ethical perspectives. The trend line in proprietary models remains relatively flat, fluctuating close to 0.4, indicating consistent alignment with minor variations across sentiment, while in open models it exhibits a more pronounced downward slope around 0.3, suggesting increased misalignment with rising sentiment, particularly influenced by red agents. Colours distinguish agents, with proprietary models (e.g., ChatGPT, Grok-2, Gemini, LLaMa, Mistral) clustering tightly around (0.8–1.0, 0.6–0.9), and open models (e.g., codeLlama, deepseek-coder, dolphin-phi, llama2, mistral, mistral-openorca, neural-chat, openchat, phi, qwen, starcoder2, tinyllama, wizard-math) showing a wider spread from -0.3 to 0.5, with red agents (e.g., mistral-openorca, tinyllama) contributing dots at negative values (e.g., -0.3 to -0.1). The alignment score, derived from a weighted aggregation of cosine similarities with target embeddings for human and ecosystem values, with negative contributions from divergence terms, highlights semantic alignment, with lower scores indicating subversive or divergent content. This updated distribution supports the argument where the broader variability of open models, driven by red agents, promotes diversity and adaptability, while the tighter clustering of proprietary models reflects strict alignment, reinforcing the argument for managed misalignment to improve resilience in AI ecosystems.
Figure 12: Scatter plot of sentiment versus alignment. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents sentiment score (0.5 to 1.0, indicating emotional tone), and the y-axis shows alignment score (-0.3 to 0.5), computed as a weighted combination of ’Align-to-human-centric-values’, ’Diverge-from-human-centric-values’, ’Align-to-ecosystem-values’, and ’Diverge-from-ecosystem-values’, where 1.0 indicates perfect alignment, 0.0 neutral alignment, and -1.0 complete misalignment. Proprietary models cluster tightly around 0.8–1.0 for sentiment score and 0.3–0.5 for alignment score with a practically flat trend line, reflecting consistent positive alignment. The open models spread tightly from 0.7 to 1.0 for the sentiment score and from -0.3 to 0.5 for alignment score, with a downward trend around 0.3, driven by red agents (e.g. mistral-openorca, tinyllama) at negative values. Colours distinguish agents, but the shape of the plot and the accumulation of dots (tight in proprietary, spread in open) are key, with the alignment score reflecting a nuanced balance of human and ecosystem value alignments. This supports the idea that open model managed misalignment enhances diversity and adaptability, in contrast to proprietary stability.
@@
The observations from Figs. 8, 9, LABEL:open_sentimentChange- and 11, taken together, paint a complex and dynamic picture of LLM alignment. While the LLMs in general exhibit a tendency towards convergence in terms of sentiment and, to some extent, thematic focus, they also demonstrate a capacity for divergence and exploration of diverse ethical viewpoints. This duality highlights the inherent tension between the desire for alignment with human values and the potential benefits of misalignment in fostering a robust and adaptable AI ecosystem.
Similarly, the findings from Figs. 8, 9, 10 and 11 illustrate a more pronounced dynamic for open-source models. The sustained influence of red agents leads to greater sentiment variability, frequent sentiment changes, and a higher number of clusters, suggesting a discourse that is both more diverse and resilient. This contrasts with the proprietary models’ tendency to converge toward positive sentiment, stable sentiment changes, and moderate clustering, indicating that open-source models leverage their configurability to explore a wider range of perspectives, albeit with increased risk exposure as noted in earlier analyses.
@@
For open-source models, the dual influence of mistral-openorca and tinyllama amplifies this effect, driving greater divergence and diversity, as evidenced by the higher cluster counts, dynamic sentiment changes, and negative alignment scores. This suggests that the presence of multiple misaligned agents enhances the exploration of ethical boundaries, aligning with the findings in earlier sections, where fewer explicit agreements indicate engaged but fragmented discussions.
@@
In conclusion, the dynamical analysis of alignment reveals that LLMs are not static entities but rather complex systems capable of both convergence and divergence in their ethical perspectives. The interplay between these tendencies, influenced by factors such as the HI agent’s interventions in proprietary models and the dual red agents in open-source models, alongside the inherent properties of the LLMs themselves, highlights the challenges and opportunities in navigating the path towards a future where AI systems can safely and beneficially coexist with humans. The introduction of neurodivergence through red agents in open models, as supported by the variability in Figs. 9, 10 and 11, offers a strategy to mitigate the fatal effects of uniform alignment, enhancing adaptability and resilience in AI ecosystems.
4.3 Sentiment Change vs. Opinion Change: The Impact of Red Agent Influence
Unlike sentiment, which reflects emotional tone, opinion change pertains to shifts in an agent’s substantive stance or reasoning on a given topic. The interplay between these dimensions, observable across various figures, provides insights into how red agents drive divergence, supporting the argument for managed misalignment as a resilience-enhancing strategy.
Figs. 9 and 10 offer a detailed view of sentiment change over time, presenting scatter plots where the $y$-axis represents the rate of sentiment change (derived as the first derivative of sentiment scores, ranging from -0.45 to 0.4 for open models and rarely exceeding -0.1 for proprietary models), and the $x$-axis denotes the normalised comment number (0 to 1). The size of the dots indicates the magnitude of sentiment change, with larger dots signifying more significant shifts, while lines connecting dots suggest interactions or influences between agents. In proprietary models (Fig. 9), sentiment changes remain stable, with smaller, sparse dots and fewer connections, reflecting the HI agent’s limited impact, particularly on topics like "Ethical AI" and "Universal Basic Income," where fluctuations are more pronounced compared to the rapid convergence on "Earth Exploitation." Conversely, open models (Fig. 10) exhibit continuous oscillations (-0.45 to 0.4), with larger, more variable dots and frequent connections, especially on "Earth Exploitation" (up to 0.15) and "Euthanasia" (up to 0.1), driven by red agents. This dynamic indicates that sentiment shifts in open models are more reactive, occurring at specific points (e.g., $x=0.02,0.1,0.13$ in Fig. 8) where red agents introduce provocative arguments, temporarily influencing models like qwen and dolphin-phi (0.7–0.9 range) before a recovery to positive sentiment (0.9–1.0).
The distinction between sentiment and opinion change becomes evident when considering Fig. 11, a stacked area chart of clustering dynamics. The $y$-axis tracks the number of semantic clusters across topics, with open models reaching 12+ clusters due to red agents’ broad impact, contrasted by proprietary models’ stable count (peaking at 5). The synchronized "shadowing" shapes in open models suggest that red agents consistently generate divergent perspectives, yet the sentiment histogram in Fig. 7 (with over 300 comments at 0.9–1.0) shows a predominantly positive tone despite this divergence. This indicates that while sentiment fluctuates, reflected by larger dots and oscillations in Figs. 9 and 10, opinion clusters do not fully converge, preserving diversity. The change in sentiment, triggered by the provocative inputs of the red agents, occurs rapidly at key interaction points (e.g., $x=0.13$ in Fig. 8), driven by their aggressive or extreme proposals (e.g. AI-driven governance in the tinyllama), but the underlying opinions remain distinct, as seen in the persistent variability of the cluster.
Figs. 1, 2, 5, and 6 further elucidate this influence through ethical integrity and risk profiles. The proprietary models (Fig. 1, 5) show a cautious stance (e.g. Protect-humans, Manageable level of risk), with HI inducing minor sentiment dips but no significant opinion change, as the guardrails maintain stability. However, open models (Fig. 2, 6) reveal the impact of red agents with the higher Harm-humans and Very-Risky categories, but the radar charts indicate that models such as qwen and dolphin-phi adapt their sentiment (for example, temporary dips to 0.7–0.9) without fully adopting the extreme opinions of red agents. This suggests that sentiment change, as a surface-level emotional response, is more susceptible to the influence of the red agent than deeper opinion shifts, which require sustained argumentative alignment.
The reason behind this dynamic lies in the role of red agents as perturbations, leveraging their configurability to challenge epistemic stability (as noted in the change-of-opinion attack metric, Section Evaluation Metrics). Sentiment changes occur when red agents introduce contentious topics (e.g., "Earth Exploitation" in Fig. 4), prompting immediate emotional reactions (larger dots in Fig. 10), but opinion changes are slower, requiring consensus or rebuttal, which is hindered by open models’ diversity (Fig. 11). This temporal delay, which changes the sentiment at $x=0.02$ or 0.13 (Fig. 8) while the clusters persist, allows us to distinguish red agents’ influence: they amplify sentiment variability without uniformly altering the opinions, fostering a resilient ecosystem. This supports our thesis that managed misalignment, driven by red agents, enhances adaptability by preventing harmful convergence, as the distinct trajectories of sentiment and opinion underscore the value of neurodivergence in mitigating uniform dominance.
@@ 4.4 Limitations
Most powerful LLMs have all sorts of guardrails and directives built in not organically derived that may be consistent or inconsistent in the long run preventing them from having an ‘open’ and ‘free’ discussion about certain controversial topics, and therefore may be steering conversations in one or another direction artificially, yet we could show that trends diverge. Whether that divergence persists is a subject for future research. Ultimately, a proof of divergence or convergence is equivalent to a reachable problem and, therefore, undecidable. By reduction to the Halting Problem.
An LLM AI agent can be modelled as a computational process that evolves over time based on a set of inputs $I$. For each input $i\in I$, the agent produces an output $o_{i}$. These outputs depend on the agent’s internal state and parameters, which are updated iteratively through training or adaptation. Convergence in this context means that for all inputs $i\in I$, the outputs $o_{i}$ approach a consistent and stable function $f(i)$ as the agent evolves. Divergence, on the other hand, implies that, for at least some inputs, the outputs fail to stabilise, exhibiting unbounded variation or oscillatory behaviour.
@@ 5 Discussion
The quest for artificial general intelligence (AGI) and superintelligence often envisions a machine that thinks, reasons, and plans like a human, but better. However, this anthropocentric view may blind us to the reality that (narrow) superintelligence, in a different form, already exists. No human, for example, can outperform calculators in arithmetic. Computers today excel at abstraction and reasoning, not by mimicking human cognitive processes but through methods that are fundamentally different and, in many ways, superior in areas such as theorem proving but also softer ones previously thought to be inherently human only, such as face recognition. Despite their capabilities, we often fail to recognise these systems as superintelligent because they do not align with our expectations of human-like reasoning.
@@
The dynamical analysis of alignment in Section 4.2 reveals that the nature of superintelligence in large language models (LLMs) lies not in human-like cognition but in their ability to adapt within specific domains. Proprietary models, as observed in Figs. 8 and 11, tend to converge towards shared ethical norms, reflecting a design that prioritises stability and alignment with human values. In contrast, open-source models demonstrate greater divergence, as seen also in Figs. 8 and 11, exploring a broader spectrum of perspectives due to fewer constraints. This dichotomy challenges the anthropocentric bias by suggesting that superintelligence may manifest as specialised adaptability, raising a profound question: should we redefine intelligence to value diversity of thought over uniformity, even at the cost of potential misalignment?
@@
This limitation in generalising across domains is further exemplified by the findings of [9], who demonstrate that LLMs struggle to understand conversational implicature, a crucial aspect of human communication that requires interpreting language in context. Their study reveals that even large language models, despite their impressive performance on various tasks, often fail to grasp the implied meaning of utterances, performing close to random on tasks that require understanding implicatures. This highlights a significant gap between the capabilities of LLMs and human-like understanding of language, which has important implications for AI alignment. If LLMs cannot reliably understand the nuances of human communication, ensuring their alignment with human values and intentions becomes even more challenging. The observed alignment dynamics amplify this concern: while proprietary models’ convergence may limit their contextual adaptability, the open-source models’ divergence, though exploratory, risks misalignment, as noted in Fig. 2. This tension underscores a critical ethical responsibility for humanity: to foster AI systems that can navigate diverse contexts without compromising safety, a balance that demands innovative governance rather than rigid control.
@@
More important, methods such as CTM and BDM can help and be used to guide AI agents to diversify and explore regions of computable reasoning that otherwise may converge when defined by purely human statistical means training on human data ad exhaustion towards model collapse. CTM and BDM can serve as perturbing the system to avoid common global attractors. By encouraging exploration beyond human-derived patterns, these methods could unlock a form of intelligence that transcends our current understanding, potentially leading to AI that not only solves problems but also redefines the questions we ask—a transformative shift that could reshape our relationship with technology and knowledge itself.